# Be like a Goldfish, Don’t Memorize! Mitigating Memorization in Generative LLMs

### v1

@@ 6.2 Adaptive Attack: Beam Search
Figure 8: Adaptive Attack: We consider an adaptive attack using beam search, which evaluates multiple sequences during generation. With a beam search width of $30$, we observe that for lower values of $k=3,4$, the $k$-GL models still outputs little to no regurgitated text. However, as $k$ increases, the adaptive attack recovers more of the original text.
@@ 7 Conclusion
We believe that goldfish loss can be a useful tool in industrial settings due to its simplicity, scalability, and relatively small impacts on model performance. While our experiments apply the loss uniformly over all documents, it can also be selectively applied during late phases of a training curriculum, or to documents from specific high-risk sources. This limits the negative impacts on utility whilst focusing mitigation where it matters most. Furthermore, in situation with plentiful but sensitive content, or low entropy text (e.g. code), one might use higher masking rates than those explored in this paper.
@@ 8 Acknowledgments
An award for computer time was provided by the U.S. Department of Energy’s (DOE) Innovative and Novel Computational Impact on Theory and Experiment (INCITE) Program. This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. Financial support was provided by the ONR MURI program and the AFOSR MURI program. Private support was provided by Capital One Bank, the Amazon Research Award program, and Open Philanthropy. Further support was provided by the National Science Foundation (IIS-2212182), and by the NSF TRAILS Institute (2229885).
@@ A.2 Hardware
Each of 1B parameter model training runs were orchestrated in Distributed Data Parallel (DDP) manner over 16 nodes of 8 GPUs. While for 7B parameter model training, we employed 4D parallelization introduced in Singh and Bhatele [2022] with 8 nodes of 8 GPUs. Each run of 1B training consumed 1280 GPU hours consuming 40 GB per GPU.
@@ Appendix B Comparison of Goldfish Loss Strategies
In Figure 10, we compare the memorization and downstream benchmark performance of goldfish loss (as introduced in Section 3) across various strategies and hyperparameter k. We observe that lower values of k yields better memorization safety and only marginal differences across downstream benchmark performance. Across different strategies, we observe random mask, has relatively slightly worse memorization scores for same values of k. This behavior is expected (Section LABEL:sec:_our-loss) since the model ends up supervising all tokens in expectations when training over multiple epochs or having duplication across batches. Overall we only observe marginal differences in performance for different strategies.
Figure 9: Validation Loss Curves: We show validation loss on RedpajamaV2 data over two models with the same hyperparameters except one was trained with standard loss and another was trained with $4$-GL with a hash width of $13$. The solid lines indicate the validation as the total number of tokens seen during training, input tokens. The dashed line represents loss in terms of the total number of tokens used for calculating the loss, supervised tokens. For the standard loss line, the dashed line is omitted as all the input tokens are used in the loss, or the input tokens are equal to the supervised tokens. We suspect that the dashed line crosses over the standard loss curve at the end because of the slower learning rates at the end.
@@ Appendix C Auxiliary Results
C.1 Membership Inference Attack

### v2

@@ 6.2 Adaptive Attack: Beam Search
Figure 8: Benchmark Performance: We pretrain 1B parameter models on 20 billion tokens as described in Section 4.1 and evaluate downstream performance on various benchmarks. We note only marginal change in performance for models trained with goldfish loss ($k=3$ and $k=4$) in comparison to the model trained with standard loss. Control refers to model trained only on RedPajama and not on wikipedia canaries.
@@ 7 Conclusion
We believe that goldfish loss can be a useful tool in industrial settings due to its simplicity, scalability, and relatively small impacts on model performance. While our experiments apply the loss uniformly over all documents, it can also be selectively applied during late phases of a training curriculum, or to documents from specific high-risk sources. This limits the negative impacts on utility whilst focusing mitigation where it matters most. Furthermore, in situation with plentiful but sensitive content, or low entropy text (e.g. code), one might use higher masking rates than those explored in this paper. We hope that goldfish loss paves the way for aiding copyright compliance rather than serving as a means to misuse private data maliciously.
@@ 8 Acknowledgments
An award for computer time was provided by the U.S. Department of Energy’s (DOE) Innovative and Novel Computational Impact on Theory and Experiment (INCITE) Program. This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. Financial support was provided by the ONR MURI program and the AFOSR MURI program. Private support was provided by Capital One Bank, the Amazon Research Award program, and Open Philanthropy. Further support was provided by the National Science Foundation (IIS-2212182), and by the NSF TRAILS Institute (2229885). We also thank the double blind reviewers for their valuable time and feedback.
@@ A.2 Hardware
Each of 1B parameter model training runs were orchestrated in Distributed Data Parallel (DDP) manner over 16 nodes of 8 GPUs. While for 7B parameter model training, we employed 4D parallelization introduced in Singh and Bhatele [2022] and Singh et al. [2024] with 8 nodes of 8 GPUs. Each run of 1B training consumed 1280 GPU hours consuming 40 GB per GPU.
@@ Appendix B Comparison of Goldfish Loss Strategies
In Figure 9, we compare the memorization and downstream benchmark performance of goldfish loss (as introduced in Section 3) across various strategies and hyperparameter k. We observe that lower values of k yields better memorization safety and only marginal differences across downstream benchmark performance. Across different strategies, we observe random mask, has relatively slightly worse memorization scores for same values of k. This behavior is expected since the model ends up supervising all tokens in expectations when training over multiple epochs or having duplication across batches. Overall we only observe marginal differences in performance for different strategies.
@@ Appendix C Auxiliary Results
Figure 10: Semantic Memorization: In addition to Rouge1 and Rouge2 measuring unigram overlap and bigram overlap, we also measure BERTScore [Zhang* et al., 2020] which is BERT embedding-based scores where a higher score suggests a closer semantic similarity to the ground truth. Despite the $4$-goldfish model’s deterrence to regenerate the exact sequences seen during training, the increased BERT embedding-based BERTScore and n-gram-based Rouge scores (in comparison to Control) suggest that paraphrases might still be leaked. This observation implies that while the model does not memorize, it still learns and retains knowledge from the underlying data.
C.1 Semantic Memorization
In the main paper, we restricted our analysis to syntactical form of memorization with metrics such as exact match rate and RougeL. As observed in Figure 1, we clearly see that goldfish loss severely restricts reproduction of training sequences verbatim. However, in this section, we aim to understand if the model preserves semantic understanding from the sequences trained with goldfish loss. Alternatively, we evaluate if the goldfish model capable of leaking paraphrased text if not exact verbatim copies.
In Figure 10, we observe that the goldfish model gets an embedding-based BERTScore of 75%, increased from the non-trained Control at 5%, and lesser than training with a standard loss at 97%. We also see a similar trend for n-gram-based Rouge scores indicating that goldfish models do generate paraphrases of training data, if not exact verbatim reproduction which is at 0% (same as Control and 85% for standard loss).
This result implies that the goldfish loss, as intended, deters the model from reproducing exact training samples during the inference phase. However, it still retains the learned knowledge from these training samples, resulting in generated text that is semantically similar to the training data without being identical.
C.2 Membership Inference Attack